OpenStudy (anonymous):

Solve the following system using the substitution method. 3x – y = –2 6x + 5y = 24 possible answers: (2/3, 4) (-2/3 , 4) (2/3 , -4) (-2/3, -4) Thank you

OpenStudy (anonymous):

3x-y=-2 y=3x+2 -> 6x+5y=24 -> 6x+5(3x+2)=24 6x+15x+10=24 21x=14 x=14/21 x=2/3 3x-y=-2 3(2/3)-y=-2 2-y=-2 y=4 the answer (2/3, 4)